# ACM SIGSOFT SEN Empirical Software Engineering: Introducing Our New Regular Column

### v1

@@ 1 Introduction
Empirical software engineering (ESE) emerged in response to the so-called “software crisis” of the 1960s, when growing system complexity revealed that theory alone was insufficient for effective software development [7]. While the term “software engineering” was first popularized during the 1968 and 1969 NATO conferences [13], it was only in the 1970s that researchers like Victor R. Basili at the University of Maryland began applying rigorous scientific methods to software engineering [3]. Basili’s pioneering work, which includes well-known frameworks such as Goal-Question-Metric [2], laid the foundations for empirical practices in the field, and guided software engineering on the path of becoming a more rigorous, evidence-based research field. Over the years, many others have provided substantial methodological contribution to ESE research, e.g., Barbara Kitchenham for systematic literature reviews [9], Carolin Seaman for qualitative studies [15], or Natalia Juristo [8] and Claes Wohlin [22] for controlled experiments, to only name a few (there are, of course, many others).
From the 80s and 90s into the 21st century, methodologies expanded from small-scale lab and student experiments to larger industrial case studies and controlled experimentation, tied together by efforts to establish reproducibility, replication, and stronger methodological robustness in software engineering research [6]. This was by no means an easy transformation, as many early SE empiricists received considerable push-back by SE researchers coming from a rationalist background [17]. But over the past two decades, the ESE discipline has matured considerably: it now includes both quantitative and qualitative approaches, leverages data mining of large software repositories, employs systematic secondary studies and mixed-method designs, places strong emphasis on open science [11], and is supported by dedicated journals and conferences such as the Empirical Software Engineering journal (EMSE) [^1] https://link.springer.com/journal/10664, the International Symposium on Empirical Software Engineering and Measurement (ESEM) [^2] https://www.esem-conferences.org, and the International Conference on Evaluation and Assessment in Software Engineering (EASE) [^3] https://conf.researchr.org/series/ease. Lastly, we also have high-quality, modern ESE books at our disposal, such as Wohlin et al. [22], Felderer and Travassos [5], or Mendez et al. [12], an evolving collection of community-driven empirical standards [14], and have arguably become much better at the methodological education of the next generation of ESE researchers.

### v2

@@ 1 Introduction
Empirical software engineering (ESE) emerged in response to the so-called “software crisis” of the 1960s, when growing system complexity revealed that theory alone was insufficient for effective software development [7]. While the term “software engineering” was first popularized during the 1968 and 1969 NATO conferences [13], it was only in the 1970s that researchers like Victor R. Basili at the University of Maryland began applying rigorous scientific methods to software engineering [3]. Basili’s pioneering work, which includes well-known frameworks such as Goal-Question-Metric [2], laid the foundations for empirical practices in the field, and guided software engineering on the path of becoming a more rigorous, evidence-based research field. Over the years, many others have provided substantial methodological contributions to ESE research, e.g., Barbara Kitchenham for systematic literature reviews [9], Carolyn Seaman for qualitative studies [15], or Natalia Juristo [8] and Claes Wohlin [22] for controlled experiments, to only name a few (there are, of course, many others).
From the 80s and 90s into the 21st century, methodologies expanded from small-scale lab and student experiments to larger industrial case studies and controlled experimentation, tied together by efforts to establish reproducibility, replication, and stronger methodological robustness in software engineering research [6]. This was by no means an easy transformation, as many early SE empiricists received considerable pushback from SE researchers coming from a rationalist background [17]. But over the past two decades, the ESE discipline has matured considerably: it now includes both quantitative and qualitative approaches, leverages data mining of large software repositories, employs systematic secondary studies and mixed-method designs, places strong emphasis on open science [11], and is supported by dedicated journals and conferences such as the Empirical Software Engineering journal (EMSE) [^1] https://link.springer.com/journal/10664, the International Symposium on Empirical Software Engineering and Measurement (ESEM) [^2] https://www.esem-conferences.org, and the International Conference on Evaluation and Assessment in Software Engineering (EASE) [^3] https://conf.researchr.org/series/ease. Lastly, we also have high-quality, modern ESE books at our disposal, such as Wohlin et al. [22], Felderer and Travassos [5], or Mendez et al. [12], an evolving collection of community-driven empirical standards [14], and have arguably become much better at the methodological education of the next generation of ESE researchers.